OpenStudy (anonymous):

integral ((e^2x)-2)/((e^2x)+3)

OpenStudy (abb0t):

Use substitution.

OpenStudy (anonymous):

Let u = (e^(2x) + 3) du = 2 (e^(2x) + 3) So this integral is of the form 1/2 du/u So the result is (1/2) ln ((e^2x) + 3)) + C

OpenStudy (anonymous):

du = 2 (e^(2x) + 3) or only du=2e^2x

hartnn (hartnn):

3 is constant, so derivative =0 so its just du = 2e^2x :)

OpenStudy (anonymous):

ok i need the final answer

hartnn (hartnn):

did you try ? what did u get ?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\ln (e^{2x}+3)-\ln (e^{2x}+3)\]

hartnn (hartnn):

nopes, that doesn't seem correct....

OpenStudy (anonymous):

yes i thought that too

hartnn (hartnn):

you split it as e^2x/ (....) - 2/ (....) right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so the first part is easy, just 1/2 ln (e^2x+3) you will need partial fractions for 2nd part...

OpenStudy (anonymous):

show me how to do this step (partial fractions)